A can of 40 marbles in it weighed 135 grams. The same can with 20 marbles weighed 75 grams. What is the mass of each marble? What is the mass of the can?
This problem was very confusing to me on my test and i just want ot know how to solve it and the answer please.

Answer by checkley71(8405)

40/(135-X)=20/(75-X) NOW CROS MULTIPLY
40(75-X)=20(135-X)
3000-40X=2700-20X
-40X+20X=2700-3000
-20X=-300
X=-300/-20
X=15 GRAMS FOR THE CANS WEIGHT.
PROOF
40/(135-15)=20/(75-15)
40/120=20/60
240=240